Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion 
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the total cost of the handling department's work in process
inventory at April 30 using the FIFO process costing method.
Answer #1
Total cost of the work in process inventory at April 30 141988
Workings:
Direct materials
To complete Work in process, April 1 11160 =18000*(1-38%)
Units started and completed 36000 =80000-44000
Work in process, April 30 12320 =44000*28%
Equivalent units 59480
Conversion
To complete Work in process, April 1 5220 =18000*(1-71%)
Units started and completed 36000 =80000-44000
Work in process, April 30 32560 =44000*74%
Equivalent units 73780
Direct materials Conversion
Costs added during April 190336 232407
Divide by Equivalent units 59480 73780
Cost per Equivalent unit 3.20 3.15
Direct materials Conversion Total
Work in process, April 30 12320 32560
X Cost per Equivalent unit 3.20 3.15
Total cost of the work in process inventory at April 30 39424 102564 141988